# Improving variational counterdiabatic driving with weighted actions and computer algebra

### v1

@@ Introduction
$\bm{\alpha}^{(\mathcal{P})}(\lambda)=\mathop{\mathrm{argmin}}_{\bm{\alpha}\in\mathbb{R}^{M}}\mathcal{S}_{\lambda}^{(\mathcal{P})}[V(\bm{\alpha})].$ (13)
While the minimization of $\mathcal{S}_{\lambda}^{(\mathcal{P})}[V]$ over any Hermitian operators gives the exact AGP independently of $\mathcal{P}(x)$, the constrained minimization generally gives a different result depending on $\mathcal{P}(x)$.
The weighted variational method extends the conventional method in the two aforementioned aspects. First, it enables us to assign weights to some of the elements $|[V-\Phi(\lambda)]_{mn}|^{2}$. This property is clearly seen by expanding $\mathcal{S}_{\lambda}^{(\mathcal{P})}[V]$ in terms of the Hamiltonian eigenbasis (see Appendix A.3 for derivation):
@@
Compared with the conventional action in Eq. (9), the deviation between $V_{mn}$ and $[\Phi(\lambda)]_{mn}$ receives an additional weight,
@@
In practical applications, the polynomial $\mathcal{P}_{\lambda}(x)$ should be designed to assign weights to important elements for specific purposes. One example is to assign weights to low-energy eigenstates to assist the adiabatic process along the ground state, which will be explored in later sections.
Second, the weighted action contains nonlocal terms. Even if the original Hamiltonian contains only local couplings, the fictitious Hamiltonian for $K\geq 2$ has nonlocal interaction terms due to $H(\lambda)^{2},H(\lambda)^{3},\dots$ appearing in $\mathcal{P}_{\lambda}(H(\lambda))$. Because of these nonlocal interactions, the weighted actions contain nonlocal terms, which enables us to effectively incorporate nonlocal information into the calculation of driving coefficients $\bm{\alpha}^{(\mathcal{P})}(\lambda)$.
In general, a higher degree polynomial $\mathcal{P}_{\lambda}(x)$ is favorable for designing a better action because we have more freedom to choose the coefficients $p_{k}(\lambda)$. However, a higher-degree polynomial requires more computation time to minimize the action (Sec. II.3). Therefore, the degree $K$ of the polynomial $\mathcal{P}_{\lambda}(x)$ should be set to the highest possible value under a practical limitation on computational time.
We remark that if the Hamiltonian has a conserved charge $J(\lambda)$ satisfying $[H(\lambda),J(\lambda)]=0$, we can make use of $J(\lambda)$ to construct more general variational characterizations and, hence, more fine-tuned weighted actions. This could be used to suppress nonadiabatic transitions more selectively, as discussed in more detail in Appendix B.2.
@@
We propose an efficient numerical procedure for solving the minimizations in Eq. (12). We first fix $\lambda$ and discuss the calculation for a single $\lambda$. Inserting the ansatz in Eq. (7) into the definition of $\mathcal{S}_{t}^{(\mathcal{P})}[V]$ in Eq. (12b), we obtain
@@
for $\mu,\nu=1,\dots,M$. The optimal $\bm{\alpha}$ is found by solving $\partial\mathcal{S}_{t}^{(\mathcal{P})}[V(\bm{\alpha})]/\partial\alpha_{\mu}=0$, which turns into a coupled linear equation,
@@
We can compute $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$ efficiently using computer algebra for spin systems with an arbitrary spin quantum number and fermionic systems. Computer algebra stores operators such as a Hamiltonian by their algebraic representations involving spin operators for spin systems and creation/annihilation operators for fermionic systems [68]. It performs computations such as addition, multiplication, and trace by directly manipulating the algebraic representations, just as humans do by hand.
@@
The computational time of $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$ with computer algebra scales polynomially in the system size for physically natural Hamiltonians, which is much more efficient than the computation with matrix representations. Consider a system on a lattice with $N$ sites. The matrix representation of an operator usually contains an exponential number of nonzero elements with respect to $N$. Thus, the computation time with matrix representations scales exponentially with the system size. In contrast, the computation time with computer algebra is roughly proportional to the number of operators involved in the computation. Assuming that the number of terms in $H(\lambda)$ is a polynomial of $N$, which is valid for almost all physical Hamiltonians, the number of operators appearing in $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$ is also a polynomial of $N$. Thus, the computational time with algebraic representations scales polynomially with the system size.
@@
For concreteness, we below focus on a general spin-$1/2$ system consisting of $N$ spins, labeled as $i=1,\dots,N$. We use $X_{i}$, $Y_{i}$, and $Z_{i}$ to denote the Pauli operators acting on the $i$th spin. The algebraic representation of an operator has the form of a sum of terms, each of which consists of a scalar coefficient and a tensor product of Pauli operators, e.g., $1.5X_{1}Y_{2}$, where we omit the identity operators acting on the remaining spins. Computer algebra manipulates this representation using algebraic relations such as $X_{j}Y_{j}=iZ_{j}$, $X_{j}X_{j}=I$, $\mathop{\mathrm{tr}}X_{j}=0$, $\mathop{\mathrm{tr}}I=2^{N}$, and so on. We develop a detailed algorithm for basic operations, such as addition, multiplication, trace, and commutators, in Appendices D.1 and D.2. These algorithms extend previous approaches [72, 73, 74, 68, 33] to make them more efficient and suitable for our purpose.
@@
Using these basic algorithms, we develop an algorithm for computing $\bm{\alpha}(\lambda)$ for a fixed $\lambda$ in Appendix D.3 and show that the computation time scales as $O(N^{K})$. The computation time is analyzed under the following four assumptions. First, $H(\lambda)$ and $A_{1},\dots,A_{M}$ are $k$-local operators with $k$ small and independent of $N$. Second, the Hamiltonian consists of $O(N)$ terms. Third, the driving operators $\{A_{\mu}\}$ satisfy either one of the following conditions: (i) the number of driving operators is $M=O(N)$, and each $A_{\mu}$ consists of $O(N^{0})$ terms; (ii) the number of driving operators is $M=O(N^{0})$, and each $A_{\mu}$ consists of $O(N)$ terms. In either case, the driving operators $\{A_{\mu}\}$ have $O(N)$ terms in total. Fourth, these $O(N)$ terms of $\{A_{\mu}\}$ are evenly distributed over $O(N)$ spins so that each spin is acted on only by $O(N^{0})$ terms. These four assumptions are satisfied by most physically natural setups, such as the Ising and Heisenberg models on a regular lattice of any dimension. Under these assumptions, the computational time obeys the $O(N^{K})$ scaling, where we neglect a multiplicative factor that mildly depends on $K$ because we will focus on a small $K$ such as $K\leq 5$ below. Due to this polynomial scaling, our algorithm can treat from tens to hundreds of spins (see Sec. IV.1 for an example of the computational time).
@@
So far, we have considered the calculation for a single value of $\lambda$. To calculate the driving coefficients for the entire range of $\lambda$, we can, of course, repeat the calculation independently for many $\lambda$’s. However, we can perform the calculation more efficiently if the Hamiltonian has the form
@@
where $F_{1},\dots,F_{\Gamma}$ are operators independent of $\lambda$, $f_{1}(\lambda),\dots,f_{\Gamma}(\lambda)$ are scalar coefficients, and $\Gamma$ is a small number such as 2 or 3. This condition on the Hamiltonian is satisfied in a large portion of the applications of variational CD driving reviewed in Sec. I. By inserting Eq. (19) into the expressions of $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$ in Eq. (17), we can separate the scalar coefficients dependent on $\lambda$ from the traces of operators independent of $\lambda$. Then, the evaluation of the traces, which is the most time consuming step, is common for all $\lambda$ and needs to be done only once (see Appendix B.3 for details). The computational complexity for these traces scales as $O(N^{K})$, similarly to the single-$\lambda$ case.
@@
We apply our general framework to ground-state evolution by choosing a polynomial $\mathcal{P}_{\lambda}(x)$ that assigns more weights to lower-energy eigenstates. This choice of a polynomial is intuitively understood by considering that we do not need to suppress nonadiabatic transitions between high-energy levels. We provide a further justification in Sec. V.1. As discussed in Sec. II.2, a larger degree polynomial $\mathcal{P}_{\lambda}(x)$ can give a better driving protocol but requires more computational time. Therefore, we specify $\mathcal{P}_{\lambda}(x)$ for each degree $K\geq 1$, allowing users to choose $K$ based on realistic limitations on computational time and desired driving quality.
We would ideally find the “best” degree-$K$ polynomial for this purpose, but this is impossible without solving the time evolution. Instead, we propose a heuristic construction of polynomials. Consider a series of polynomials,
$\mathcal{P}_{\lambda}^{\mathrm{GS},K}(x)=(x-E_{\lambda}^{(K)})^{K}$ (20)
for $K=1,2,\dots$, where GS stands for “ground state,” and $E_{\lambda}^{(K)}$ is a constant determined later so that the weights are concentrated on low-energy states as much as possible. This polynomial $\mathcal{P}_{\lambda}^{\mathrm{GS},K}(x)$ corresponds to choosing the coefficients $p_{k}(\lambda)\coloneqq\binom{K}{k}(E_{\lambda}^{(K)})^{K-k}$ in Eq. (10), where $\binom{x}{y}$ denotes the binomial coefficient. Below, we use the notation $\mathcal{S}_{\lambda}^{(K)}[V]\equiv\mathcal{S}_{\lambda}^{(\mathcal{P}^{\mathrm{GS},K})}[V]$, $\bm{\alpha}^{(K)}(\lambda)\equiv\bm{\alpha}^{(\mathcal{P}^{\mathrm{GS},K})}(\lambda)$, $Q_{\mu\nu}^{(K)}(\lambda)\equiv Q_{\mu\nu}^{(\mathcal{P}^{\mathrm{GS},K})}(\lambda)$, $r_{\mu}^{(K)}(\lambda)\equiv r_{\mu}^{(\mathcal{P}^{\mathrm{GS},K})}(\lambda)$, and $w_{mn}^{(K)}\equiv w_{mn}^{(\mathcal{P}^{\mathrm{GS},K})}(\lambda)$ for conciseness. In this notation, the minimization problem for determining the coefficient reads
@@
The explicit algebraic and expanded forms of the action $\mathcal{S}_{\lambda}^{(K)}[V]$ are
$\mathcal{S}_{\lambda}^{(K)}[V]$
$=\bigl{\|}\partial^{\prime}_{\lambda}\{[{H}(\lambda)-E_{\lambda}^{(K)}]^{K}\}-i\hbar^{-1}[[{H}(\lambda)-E_{\lambda}^{(K)}]^{K},V]\bigr{\|}^{2}$ (22a)
$=\frac{1}{\hbar^{2}}\sum_{nm}\biggl{(}\Bigl{\{}[\epsilon_{m}(\lambda)-E_{\lambda}^{(K)}]^{K}-[\epsilon_{n}(\lambda)-E_{\lambda}^{(K)}]^{K}\Bigr{\}}^{2}$
×$\hskip 80.00012pt\times\bigl{|}[V-\Phi(\lambda)]_{mn}\bigr{|}^{2}\biggr{)}+\mathrm{const.}\quad$ (22b)
The degree-one ($K=1$) action equals the conventional action in Eq. (6b), which justifies the notation $\mathcal{S}_{\lambda}^{(1)}[V]$ for the conventional action.
We analyze this weighted action by examining the additional weight $w_{mn}^{(K)}(\lambda)$ onto the $(m,n)$th element, introduced in Eq. (15). The weight is explicitly calculated as (see Appendix A.4 for derivation)
$w_{mn}^{(K)}(\lambda)=\!\sum_{s=-(K-1)}^{K-1}\biggl{\{}(K-|s|)$
×$\qquad\quad\times[\epsilon_{n}(\lambda)-E_{\lambda}^{(K)}]^{K-1-s}[\epsilon_{m}(\lambda)-E_{\lambda}^{(K)}]^{K-1+s}\biggr{\}}.$ (23)
@@
Moreover, if $[\epsilon_{n}(\lambda)-E_{\lambda}^{(K)}]$ and $[\epsilon_{m}(\lambda)-E_{\lambda}^{(K)}]$ have the same sign, $w_{mn}^{(K)}(\lambda)$ lies between $w_{n}^{(K)}(\lambda)$ and $w_{m}^{(K)}(\lambda)$. If we further assume $w_{n}^{(K)}(\lambda)\simeq w_{m}^{(K)}(\lambda)$, this fact implies that the weight is approximated as $w_{mn}^{(K)}(\lambda)\simeq w_{n}^{(K)}(\lambda)\simeq w_{m}^{(K)}(\lambda)$ (see Appendix A.4 for the derivation of these properties). Due to these properties, we regard $w_{n}^{(K)}(\lambda)$ as the weight associated with the $n$th eigenstate for heuristic discussions in the following.
@@
The function $\Omega_{\lambda}^{(K)}(E)$ is the average energy over the weight $[\epsilon_{n}(\lambda)-E]^{2K-2}$, and a smaller $\Omega_{\lambda}^{(K)}(E)$ implies that the weight $[\epsilon_{n}(\lambda)-E]^{2K-2}$ is more concentrated on low-energy states. Therefore, by choosing $E_{\lambda}^{(K)}$ as the minimizer of $\Omega_{\lambda}^{(K)}(E)$, we can make the resulting weight $w_{n}^{(K)}(\lambda)=K^{2}[\epsilon_{n}(\lambda)-E_{\lambda}^{(K)}]^{2K-2}$ emphasize low-energy states as much as possible. We discuss the typical behavior of the function $\Omega_{\lambda}^{(K)}(E)$ and the optimal energy shift $E_{\lambda}^{(K)}$ in Appendix C.1.
@@
The numerical implementation of the weighted variational method for ground-state evolution involves two optimization problems, one for the action $\mathcal{S}_{\lambda}^{(K)}[V(\bm{\alpha})]$ and the other for $\Omega_{\lambda}^{(K)}(E)$. The minimization of $\mathcal{S}_{\lambda}^{(K)}[V(\bm{\alpha})]$ is efficiently performed with computer algebra, as discussed in the general framework in Sec. II.3. The minimization of $\Omega_{\lambda}^{(K)}(E)$ is also efficiently computed using computer algebra. To do so, we rewrite the function $\Omega_{\lambda}^{(K)}(E)$ without explicitly using the energy eigenvalues as
@@
where we define $\omega_{k}(\lambda)\coloneqq\mathop{\mathrm{tr}}[H(\lambda){}^{k}]$ for $k=0,\dots,2K-1$. To minimize this function, we first compute the values of $\omega_{k}(\lambda)$ using computer algebra. Then, Eq. (26) becomes a rational function of $E$ with known coefficients. The minimization of $\Omega_{\lambda}^{(K)}(E)$ is done by deriving the analytical expression of $d\Omega_{\lambda}^{(K)}/dE$ and solving $d\Omega_{\lambda}^{(K)}/dE=0$ using standard numerical methods such as the Newton method. See Appendix C.1 for a further discussion on this optimization.
We summarize the entire algorithm for the ground-state evolution in Appendix D.4, where we show that the overall computational complexity scales as $O(N^{K})$, similarly to the general framework. Furthermore, if the $\lambda$ dependence of the Hamiltonian is in the form of Eq. (19), we can reduce the cost of evaluating the trace $\omega_{k}(\lambda)$ by separating the $\lambda$-dependent coefficients from the $\lambda$-independent traces, as was done for $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$ in the general framework (see Appendix C.2 for details).
@@
We test the weighted variational method applied to ground-state evolution with a quantum annealing protocol on the transverse-field Ising model with random couplings,

### v3

@@ Introduction
$\bm{\alpha}^{(\mathcal{P})}(\lambda)=\mathop{\mathrm{argmin}}_{\bm{\alpha}\in\mathbb{R}^{M}}\mathcal{S}_{\lambda}^{(\mathcal{P})}[V(\bm{\alpha})],$ (13)
where $V(\bm{\alpha})$ denotes the ansatz in Eq. (7). While the minimization of $\mathcal{S}_{\lambda}^{(\mathcal{P})}[V]$ over any Hermitian operators gives the exact AGP irrespectively of $\mathcal{P}_{\lambda}(x)$, the constrained minimization generally gives different results depending on $\mathcal{P}_{\lambda}(x)$.
The weighted variational method extends the conventional method in the two aforementioned aspects. First, it enables us to assign weights to some of the matrix elements $|[V-\Phi(\lambda)]_{mn}|^{2}$. This property is clearly seen by expanding $\mathcal{S}_{\lambda}^{(\mathcal{P})}[V]$ in terms of the Hamiltonian eigenbasis (see Appendix A.2 for derivation):
@@
Compared with the conventional action in Eq. (9), the deviation between $V_{mn}$ and $\Phi(\lambda)]_{mn}$ receives an additional weight,
@@
In a practical application, the polynomial $\mathcal{P}_{\lambda}(x)$ should be designed to assign weights to important elements for a specific purpose. One example is to assign weights to low-energy eigenstates to assist the adiabatic process along the ground state, which will be explored in later sections.
Second, the weighted action contains nonlocal terms. Even if the original Hamiltonian contains only local couplings, the fictitious Hamiltonian for $K\geq 2$ has nonlocal interaction terms due to $H(\lambda)^{2},H(\lambda)^{3},\dots$ appearing in $\mathcal{P}_{\lambda}(H(\lambda))$. Due to these nonlocal interactions, the weighted actions contain nonlocal terms, enabling us to effectively incorporate nonlocal information into the driving coefficients $\bm{\alpha}^{(\mathcal{P})}(\lambda)$.
The comparison between the conventional and weighted variational methods is summarized more intuitively in Fig. 1(b). The “hammer”, i.e., the driving term $V$, from the weighted variational method is tuned to work best for a specific part of the energy spectrum, such as the low-energy state. This is in contrast to the hammer from the conventional method, which is not optimized for the specific part of the spectrum. By this tuning, the hammer is expected to hit the target state more accurately, thereby suppressing deviations from the target dynamics more effectively.
In general, a higher degree polynomial $\mathcal{P}_{\lambda}(x)$ is favorable for designing a better action because we have more freedom to choose the coefficients $p_{k}(\lambda)$. However, a higher-degree polynomial requires more computation time to minimize the action (Sec. II.4). Therefore, the degree $K$ of the polynomial $\mathcal{P}_{\lambda}(x)$ should be set to the highest possible number under a practical limitation on computational time.
We remark that if the Hamiltonian has a conserved charge $J(\lambda)$ satisfying $H(\lambda),J(\lambda)]=0$, we can make use of $J(\lambda)$ to construct more general variational characterizations and, hence, more fine-tuned weighted actions. This could be used to suppress nonadiabatic transitions more selectively, as discussed in more detail in Appendix A.3.
@@
The minimization in Eq. (12) is efficiently solved by an algebraic method, avoiding the costly manipulation of matrix expressions and the diagonalization. This is enabled because the algebraic expression of $\mathcal{S}^{(\mathcal{P})}_{\lambda}[V]$ in Eq. (12b) involves no explicit matrix elements.
We first fix $\lambda$ and discuss the calculation for a single $\lambda$. Inserting the ansatz in Eq. (7) into the definition of $\mathcal{S}_{\lambda}^{(\mathcal{P})}[V]$ in Eq. (12b), we obtain
@@
for $\mu,\nu=1,\dots,M$. The optimal $\bm{\alpha}$ is found by solving $\partial\mathcal{S}_{\lambda}^{(\mathcal{P})}[V(\bm{\alpha})]/\partial\alpha_{\mu}=0$, which turns into a coupled linear equation,
@@
We can compute $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$ efficiently using computer algebra for spin systems with an arbitrary spin quantum number and fermionic systems. Computer algebra stores operators, such as a Hamiltonian, by their algebraic representations composed of elementary operators, such as spin operators for spin systems and creation/annihilation operators for fermionic systems [72]. It performs computations such as addition, multiplication, and trace by directly manipulating the algebraic representations, just as humans do by hand.
@@
In physically natural setups, the computational time of $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$ with computer algebra scales polynomially in system size, which is much more efficient than the computation with matrix representations. Consider a system on a lattice with $N$ sites. The matrix representation of an operator usually contains an exponential number of nonzero elements with respect to $N$. Thus, the computation time with matrix representations scales exponentially with the system size. In contrast, the computation time with computer algebra is roughly proportional to the number of elementary operators involved in the computation. Assuming that the number of elementary operators in $H(\lambda)$ and $\{A_{\mu}\}$ is a polynomial of $N$, which is valid in almost all physically natural setups, the number of elementary operators appearing in $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$ is also a polynomial of $N$. Thus, the computational time with algebraic representations scales polynomially with the system size.
@@
For concreteness, we below focus on a general spin-$1/2$ system consisting of $N$ spins, labeled by $i=1,\dots,N$. We use $X_{i}$, $Y_{i}$, and $Z_{i}$ to denote the Pauli operators acting on the $i$th spin. The algebraic representation of an operator has the form of a sum of terms, and each term consists of a scalar coefficient and a tensor product of Pauli operators, e.g., $1.5X_{1}Y_{2}$, where we omit the identity operators acting on the remaining spins. Computer algebra manipulates this representation using algebraic relations such as $X_{j}Y_{j}=iZ_{j}$, $X_{j}X_{j}=I$, $\mathop{\mathrm{tr}}X_{j}=0$, $\mathop{\mathrm{tr}}I=2^{N}$, and so on. We develop a detailed algorithm for basic operations such as addition, multiplication, trace, and commutators in Appendices C.1 and C.2. These algorithms inherit basic ideas from existing computer-algebra frameworks [76, 77, 78, 72, 33, 79] and make them more efficient and suitable for our purposes. Using these basic algorithms, we present a concrete algorithm for computing $\bm{\alpha}^{(\mathcal{P})}(\lambda)$ for a fixed $\lambda$ in Appendix C.3.
@@
Our computer algebra algorithm for general spin-$1/2$ systems can compute $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$ in $O(N^{K})$ time, as shown in Appendix C.3. This computation time is analyzed under the following four assumptions. First, $H(\lambda)$ and $A_{1},\dots,A_{M}$ are $k$-local operators with $k$ small and independent of $N$. Second, the Hamiltonian consists of $O(N)$ terms. Third, the driving operators $\{A_{\mu}\}$ satisfy either one of the following conditions: (i) the number of driving operators is $M=O(N)$, and each $A_{\mu}$ consists of $O(N^{0})$ terms; (ii) the number of driving operators is $M=O(N^{0})$, and each $A_{\mu}$ consists of $O(N)$ terms. In either case, the driving operators $\{A_{\mu}\}$ have $O(N)$ terms in total. Fourth, these $O(N)$ terms of $\{A_{\mu}\}$ are evenly distributed over $O(N)$ spins so that each spin is acted on only by $O(N^{0})$ terms. These four assumptions are satisfied by most physically natural setups, such as local control of the Ising and Heisenberg models on a regular lattice of any dimension. Under these assumptions, the computational time scales as $O(N^{K})$, where we neglect a multiplicative factor that mildly depends on $K$ because we will focus on a small $K$ such as $K\leq 5$ below. Due to this polynomial scaling, our algorithm can treat from tens to hundreds of spins (see Sec. IV.1 for an example of the computational time).
This $O(N^{K})$-time algebraic computation is followed by a standard numerical method to solve the linear equation in Eq. (18). This step takes $O(M^{3})$ time if we use exact methods, such as the Householder QR factorization [80]. Alternatively, iterative (approximate) methods, such as the conjugate gradient method, can solve Eq. (18) in $O(RM^{2})$ time, where $R$ is the number of iterations required for convergence within a specified tolerance [80].
@@
So far, we have considered the calculation for a single value of $\lambda$. To calculate the driving coefficients for the entire range of $\lambda$, we can, of course, repeat the calculation independently for many values of $\lambda$. However, we can perform the calculation more efficiently if the Hamiltonian has the form
@@
where $F_{1},\dots,F_{\Gamma}$ are operators independent of $\lambda$, $f_{1}(\lambda),\dots,f_{\Gamma}(\lambda)$ are scalar coefficients, and $\Gamma$ is a small number such as 2 or 3. This condition on the Hamiltonian is satisfied in a large portion of the applications of variational CD driving reviewed in Sec. I. By inserting Eq. (19) into the expressions of $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$ in Eq. (17), we can separate the scalar coefficients dependent on $\lambda$ from the traces of operators independent of $\lambda$. Then, the evaluation of the traces, which is the most time consuming step, is common for all $\lambda$ and needs to be done only once (see Appendix A.4 for details). With our concrete algorithm for spin-$1/2$ systems, the computational complexity for these traces scales as $O(N^{K})$, similarly to the single-$\lambda$ case (see Appendix C.3).
@@
To assist ground-state evolution, we choose a polynomial $\mathcal{P}_{\lambda}(x)$ that assigns more weights to lower-energy eigenstates. A simple reasoning behind this choice is that we do not need to suppress nonadiabatic transitions between high-energy levels. We provide a further justification in Sec. V.1. As discussed in Sec. II.3, a larger-degree polynomial $\mathcal{P}_{\lambda}(x)$ can give a better driving protocol but requires more computational time. Therefore, we specify $\mathcal{P}_{\lambda}(x)$ for each degree $K\geq 1$, allowing users to choose $K$ based on realistic limitations on computational time. For every $K\geq 1$, we heuristically take a polynomial of the form,
$\mathcal{P}_{\lambda}^{\mathrm{GS},K}(x)\coloneqq(x-E_{\lambda}^{(K)})^{K},$ (20)
where GS stands for “ground state,” and $E_{\lambda}^{(K)}$ is a constant optimized later so that the weights are concentrated on low-energy states as much as possible. This polynomial $\mathcal{P}_{\lambda}^{\mathrm{GS},K}(x)$ corresponds to choosing the coefficients $p_{k}(\lambda)=\binom{K}{k}(-E_{\lambda}^{(K)})^{K-k}$ in Eq. (10), where $\binom{x}{y}$ denotes the binomial coefficient. Below, we use the notation $\mathcal{S}_{\lambda}^{(K)}[V]\equiv\mathcal{S}_{\lambda}^{(\mathcal{P}^{\mathrm{GS},K})}[V]$, $\bm{\alpha}^{(K)}(\lambda)\equiv\bm{\alpha}^{(\mathcal{P}^{\mathrm{GS},K})}(\lambda)$, $Q_{\mu\nu}^{(K)}(\lambda)\equiv Q_{\mu\nu}^{(\mathcal{P}^{\mathrm{GS},K})}(\lambda)$, $r_{\mu}^{(K)}(\lambda)\equiv r_{\mu}^{(\mathcal{P}^{\mathrm{GS},K})}(\lambda)$, and $w_{mn}^{(K)}\equiv w_{mn}^{(\mathcal{P}^{\mathrm{GS},K})}(\lambda)$ for conciseness. In this notation, the minimization problem for determining the coefficient reads
@@
The degree-one ($K=1$) action equals the conventional action in Eq. (6b), consistent with the notation $\mathcal{S}_{\lambda}^{(1)}[V]$ for the conventional action.
We analyze the weighted action $\mathcal{S}_{\lambda}^{(K)}[V]$ by examining the additional weight $w_{mn}^{(K)}(\lambda)$ onto the $(m,n)$th element, introduced in Eq. (15). The weight is explicitly calculated as (see Appendix B.1 for derivation)
$w_{mn}^{(K)}(\lambda)$
$=\frac{\bigl\{[\epsilon_{m}(\lambda)-E_{\lambda}^{(K)}]^{K}-[\epsilon_{n}(\lambda)-E_{\lambda}^{(K)}]^{K}\bigr\}^{2}}{[\epsilon_{m}(\lambda)-\epsilon_{n}(\lambda)]^{2}}$
$=\!\sum_{\smash{s=-(K-1)}}^{K-1}\!(K-|s|)[\epsilon_{n}(\lambda)-E_{\lambda}^{(K)}]^{K-1-s}[\epsilon_{m}(\lambda)-E_{\lambda}^{(K)}]^{K-1+s}.$ (22)
@@
Moreover, when $\epsilon_{n}(\lambda)-E_{\lambda}^{(K)}]$ and $\epsilon_{m}(\lambda)-E_{\lambda}^{(K)}]$ have the same sign, $w_{mn}^{(K)}(\lambda)$ lies between $w_{n}^{(K)}(\lambda)$ and $w_{m}^{(K)}(\lambda)$. This fact implies that the weight is approximated as $w_{mn}^{(K)}(\lambda)\simeq w_{n}^{(K)}(\lambda)\simeq w_{m}^{(K)}(\lambda)$ when the state-dependent weights further satisfy $w_{n}^{(K)}(\lambda)\simeq w_{m}^{(K)}(\lambda)$ (see Appendix B.1 for the derivation of these properties). Due to these properties, we regard $w_{n}^{(K)}(\lambda)$ as the weight associated with the $n$th eigenstate for heuristic discussions in the following.
@@
The function $\Omega_{\lambda}^{(K)}(E)$ is the average energy over the weight $\epsilon_{n}(\lambda)-E]^{2K-2}$, and a smaller $\Omega_{\lambda}^{(K)}(E)$ implies that the weight $\epsilon_{n}(\lambda)-E]^{2K-2}$ is more concentrated on low-energy states. Therefore, by choosing $E_{\lambda}^{(K)}$ as the minimizer of $\Omega_{\lambda}^{(K)}(E)$, we can make the resulting weight $w_{n}^{(K)}(\lambda)=K^{2}[\epsilon_{n}(\lambda)-E_{\lambda}^{(K)}]^{2K-2}$ emphasize low-energy states as much as possible. We discuss the typical behavior of the function $\Omega_{\lambda}^{(K)}(E)$ and the energy shift $E_{\lambda}^{(K)}$ in Appendix B.2.
We remark that the particular polynomial form in Eq. (20) is not the only possible choice. For example, we could directly use the general polynomial form in Eq. (10) and determine the coefficients $p_{1}(\lambda),\dots,p_{K}(\lambda)$ by a multivariate numerical optimization similar to Eq. (24) so that the corresponding weight concentrates the most in the lower-energy state. Nevertheless, the particular form in Eq. (20) is simpler and performs sufficiently well, as demonstrated in the following. Furthermore, the particular form is expected to be robust, as its behavior can be theoretically analyzed and supported (Sec. V.1).
@@
The numerical implementation of the weighted variational method for ground-state evolution involves two optimization problems, one for the action $\mathcal{S}_{\lambda}^{(K)}[V(\bm{\alpha})]$ and the other for $\Omega_{\lambda}^{(K)}(E)$. The minimization of $\mathcal{S}_{\lambda}^{(K)}[V(\bm{\alpha})]$ is efficiently performed with computer algebra, as discussed in the general framework in Sec. II.4. The minimization of $\Omega_{\lambda}^{(K)}(E)$ is also efficiently computed using computer algebra. To do so, we rewrite the function $\Omega_{\lambda}^{(K)}(E)$ without explicitly using the energy eigenvalues as
@@
where we define $\omega_{k}(\lambda)\coloneqq\mathop{\mathrm{tr}}[H(\lambda){}^{k}$] for $k=0,\dots,2K-1$. To minimize this function, we first compute the values of $\omega_{k}(\lambda)$ using computer algebra, which is completed in polynomial time in system size. Then, Eq. (25) becomes a rational function of $E$ with known coefficients, and the minimization of $\Omega_{\lambda}^{(K)}(E)$ is easily conducted by standard numerical methods such as the Newton method. See Appendix B.2 for further discussion on this optimization.
For general spin-$1/2$ systems, we summarize the entire algorithm for the ground-state evolution in Appendix C.4, where we show that the time for the algebraic computation scales as $O(N^{K})$, similarly to the general framework. Furthermore, if the $\lambda$ dependence of the Hamiltonian is in the form of Eq. (19), we can reduce the cost of evaluating the trace $\omega_{k}(\lambda)$ by separating the $\lambda$-dependent coefficients from the $\lambda$-independent traces, as has been done for $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$ in the general framework, without changing the overall $O(N^{K})$ scaling (see Appendices B.3 and C.4 for details).
@@
We test the weighted variational method with a quantum annealing protocol on a transverse-field Ising model with inhomogeneous parameters. Quantum annealing solves optimization problems via ground-state evolution, with a variety of real-world applications proposed, ranging from chemistry to finance [81]. An optimization problem is encoded in the final Hamiltonian so that its ground state corresponds to the optimal solution. The solution is found by adiabatically reaching the ground state and performing a measurement on it. Since the target ground state is unknown, the control protocol must be determined without the target information. To meet this requirement, variational CD driving has been used to enhance quantum annealing protocols [28, 29, 30, 31, 32, 33, 43, 44, 45, 46, 47, 48, 49]. The conventional method increases the final fidelity, i.e., the probability of obtaining the correct optimal solution, but the fidelity drops considerably for large systems [29].
Quantum annealing is precisely one of the situations that suffer from the two shortcomings of the conventional variational method (Sec. II.3). First, the desired time evolution is along the unknown ground state. Thus, suppressing nonadiabatic transitions between high-energy states is irrelevant. Second, the local magnetic fields and coupling constants between spins are spatially inhomogeneous to encode a specific problem, and the solution to the problem is determined by globally integrating all information of these system parameters. Thus, local driving coefficients determined solely from the local parameters are unfavorable. Our proposal, which overcomes these two shortcomings, can indeed enhance the fidelity of quantum annealing processes, as we demonstrate below.
Consider the ground-state evolution of the following transverse-field Ising model with random couplings,